OpenStudy (anonymous):

The population P of a city can be modeled by P=240,360^0.012r where t=0 represents 2000. What is the population in the year 2000?

OpenStudy (anonymous):

P(t) = 240360*e^ 0.012t e^0.012t = P(t)/ 240360 0.012*t = ln(P(t)/ 240360) t = (1/0.012)*ln(P(t)/ 240360) If P(t)=275000 (time to use a calculator) t = 11.22 t=0 is start of 2000, so the population will hit 275000 11.22 years later, ie in about mid-March 2011 <<<

OpenStudy (michele_laino):

@khuckelbery Sorry, if t=0 then P=240,360, I think

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

no prob